OpenStudy (anonymous):

simplify this radical expression or simplify this quotient. i will give the equation below

OpenStudy (anonymous):

\[\frac{ 5 }{ \sqrt{2}-1 }\]

OpenStudy (anonymous):

First of all welcome to openstudy:) I'm not sure if you already learned this since it is not right to have a radical expression at the denominator you'll have to multiply this fraction by it's conjugate \(\LARGE (\frac{5}{\sqrt{2}-1})(\frac{\sqrt{2}+1}{\sqrt{2}+1})\) simplify it and tell me what you get

OpenStudy (anonymous):

actually i did learn it but i got confused after that part and thanks for the welcoming

OpenStudy (anonymous):

\[\left(\begin{matrix}5 \sqrt{2}+5\ /2-1\end{matrix}\right)\]

OpenStudy (anonymous):

its suppose to be in fraction form sorry

OpenStudy (anonymous):

HELLO??

OpenStudy (anonymous):

\[\frac{ 5\sqrt{2}+5 }{ 1 }\]

OpenStudy (anonymous):

thats what i got do i simplify anymore???

OpenStudy (anonymous):

will it be \[5\sqrt{2}+5\]

OpenStudy (anonymous):

sorry i was away for my dinner you already got it right! :) you can just ignore the denominator since the value is 1 anything that is divided by 1 is simply the valus of the numerator so it will be \(5\sqrt{2}+5\ or\ (5(\sqrt{2}+1)\)

OpenStudy (anonymous):

yeah ;p

OpenStudy (anonymous):

its ok thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):



OpenStudy (anonymous):

lol :) if you need more help just ask us

OpenStudy (anonymous):

ok